3. () Use only NAND gates to implement the Boolean function F AC +BC. (ii) Use only NOR gates to implement the Boolean function F AB+BC. Write the truth tables and draw the logic circuits for the following Boolean functions: (i) F A +BC. (ii) F AB +C+D. 4.
Answer #1

Please comment below, If you find any difficulty in understanding

Please Upvote, If you are impressed with my answer

3.

i) Implementing F = AC ' + BC using only NAND Gates(3) Using only NAND Gates Tmplementation: (AC (BC) By Demorgans Laii) Implementing F = AB + BC using only NOR Gatesoji) Using only OR Gates fA 1B

4. Truth tables and Logic circuits for

i) F = A + BC '4) togic citcuit Tenuth table

ii) F = AB + C ' + DC. Tuth table o O